OpenStudy (anonymous):

A highway is shown as a straight line on a coordinate grid. One end of the highway is located at (5, 9) and the other end is located at (3, -7). Sam performed the steps shown to find the coordinates of the midpoint of the highway.

OpenStudy (anonymous):

Use the midpoint formula. http://en.wikipedia.org/wiki/Midpoint

OpenStudy (anonymous):

Step 1. x1 = 5, y1 = 3, x2 = 9, y2 = -7 Step 2. x= (5+9) y=(3-7) 2 2 Step 3. x=14 y=-4 2 2 Step 4. Step 5. x = 7, y = -2 In which step did Sam first make an error?

OpenStudy (anonymous):

never mind i got it, it was step number two because he used an incorrect formula

OpenStudy (anonymous):

I was also going to say Step 1 because he wrote the coordinates wrong into his formula.

OpenStudy (anonymous):

oh yea your right i just noticed lol, Thanks!

OpenStudy (anonymous):

You're very welcome.